[Brahms-l] Shift report 20040308 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Mon Mar 08 2004 - 16:03:58 EST
------------------------------------------------------------
Date: Mon Mar  8 16:03:58 2004, Shift: 08:00-16:00
Supervisor: JHL, with: 
Log book#15 pages 79 - 79

Shift Summary:
------------------------------------------------------------
1. Beam for ~3hrs. 
   One store (56 bunches) from the previous shift, and
   another short one (45 bunches) followed by RF problem

2. Data:
   Finished MRS 6B 4M Trig.3 
   Started MRS 6A 4M Trig.3
   FS 8A Full Field 1.3M to go.

   ZDC test for Low Energy running 
   (FV: See e-log for details)
3. No Detector Problems 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 8 16:04:21 2004

This archive was generated by hypermail 2.1.8 : Mon Mar 08 2004 - 16:04:41 EST